What percentage of the mass of the isotope boron-10 do its neutrons contribute?

proton = 1.0073 amu

neutron = 1.0087 amu

electron = 5.4858×10-4 amu

............%


1
Expert's answer
2021-01-27T04:35:54-0500

5 neutron 5 protons and 5 electrons.

neutron = 5×1.0087 = 5.0435

total = 5×1.0073 + 5×1.0087 + 5×5.4858x10-4

total = 5.0365 + 5.0435 + 0.0027429

total = 10.0827429

percentage of neutron = (neutron/total )x100

=[ (5.0435)/(10.0827429) ] × 100

= 50.02111082

= 50.02